OpenStudy (anonymous):

what is the factor of x2-6x-8

OpenStudy (lgbasallote):

are you sure the last nummber is 08?

OpenStudy (lgbasallote):

-8 *

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

i doubt it's factorable then

OpenStudy (anonymous):

i dont know

OpenStudy (lgbasallote):

are you sure @Wired ?

OpenStudy (anonymous):

4 times 2

OpenStudy (anonymous):

4+2=6 and 4x2=8

OpenStudy (lgbasallote):

if you have written this correctly @elainediaz then it is not factorable because the only factors of 8 that add up to -6 is -4 x -2 and that results to POSITIVE 8 not negative

OpenStudy (anonymous):

Bah, same damn typo twice lol. \[(x+a)(x-b) = x^{2}+ax-bx-ab = x^{2}+(a-b)x-ab\] Yeah, not factorable.

OpenStudy (anonymous):

ok its not factorable

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i di positive

OpenStudy (anonymous):

This would be consistent with finding a complex root. Use the quadratic formula to find the root values of \[x ^{2} - 6x - 8; a = 1, b = -6, c = -8\]

OpenStudy (anonymous):

Quadratic formula \[- b \pm \sqrt{b^2 -4ac}\div2a\]

OpenStudy (anonymous):

@jenxcc question is to factorize and not to find the values of x..

OpenStudy (anonymous):

Factorizing this cannot be factored as a real root, true, but can be as a complex root. So, \[(x-(3+\sqrt{17})(x-(3-\sqrt{17})\]

OpenStudy (theviper):

\[\LARGE{\color{red}{\frac{-b \sqrt{b^2-4ac}}{2a}}}\]

OpenStudy (theviper):

use this & get ur ans:)